I need help with this problem... The manager of a bulk foods establishment sells a trail mix for $8 per pound and the premium cashews for $15 per pound. The manager wishes to make a 105-pound trail mix cashew mixture that will sell for $12 per pound. How many pounds of each should be used?
Answer by rwm(914)

t+c=105
8t+15c=105*12
c=60 t=45